If A and B are two matrices such that B= -A-1 BA, then (A+B)2= ?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Ishita
 
B =  –A–1 BA
 
hence
 
AB= –BA
 
now
 
(A+B)2 = (A+ B)(A + B) = A2 + B2 + AB + BA
 
 = A2 + B2 +AB – AB
 = A2 + B2 
 
 
Regards
Arun (askIITians forum expert)